OpenStudy (anonymous):

Two end points of a line segment are (2, 8) and (-4, -8). What are the coordinates of the point on the line through which its bisector passes? (5, -6) (3, 8) (-2, 0) (-1, 0)

OpenStudy (anonymous):

A bisector cuts a segment in half, correct? so were basically finding the midpoint of the segment. when you have two points (x1,y1) (x2,y2) the coordinate of the midpoint is ((x1+x2)/2, (y1+y2)/2)) can you plug the numbers in?

OpenStudy (anonymous):

-2,0 right

OpenStudy (anonymous):

Well lets see. (2+(-4))/2=-2/2=-1 (8+(-8))/2=0/2=0 (-1,0)

OpenStudy (anonymous):

oh ok i see where i made my mistake. Thank you I understand now. Can you help me with a couple more